If three cards are randomly drawn from a standard deck how many ways are there to have
A.)at least one black?
b.) All the same color?

Answer by stanbon(75887)

If three cards are randomly drawn from a standard deck how many ways are there to have
A.)at least one black?
-------
# of ways to succeed = total # of triples - # of triples with 3 red cards
= 52C3-26C3 = 19500
----
# of random triples = 52C3 = 22100
-----
P(triple with at least one black) = 19500/22100 = 0.8824
----------------
b.) All the same color?
# of ways to succeed:: 26C3 + 26C3 = 5200
---
P(all same color) = 5200/52C3 = 0.2353
=========================
Cheers,
Stan H.
========================